I have 40 bills in my cash drawer. Some are $5 bills. The rest are $10 bills.
The total amount of the 40 bills is $260.
How many $5 bills and how many $10 bills are in my cash drawer ???

Answer by oberobic(2304)

Like coin problems, you have to keep track of the count of the bills and their value.
x = number of $5 bills
5x = value of the $5 bills
y = number of $10 bills
10y = value of the $10 bills
.
x + y = 40
so
x = 40-y
.
5x + 10y = 260
.
substitute for x
.
5(40-y) + 10y = 260
200 -5y + 10y = 260
5y = 60
y = 12
.
x = 40-y
x = 28
.
Check the value to determine if this is the correct answer.
.
5*28 = 140
10 * 12 = 120
140 + 120 = 260
Correct.
.
Answer: You have 28 $5 bills and 12 $10 bills.
.
Done.